OpenStudy (anonymous):

et f=, evaluate the line integral below where c is the parabola on the xy plane: y =x^2, 0<=x<=1

OpenStudy (anonymous):

are you a teacher or how do you know all this so well?

OpenStudy (rational):

line integral along curve \(C\) : \[\int_C x dx + (2x+3y)dy\]

OpenStudy (rational):

parameterize and evaluate

OpenStudy (rational):

given curve : \(y=x^2\) see if below parameterization works : \(x = t\) \(y = t^2\)

OpenStudy (rational):

yes it does \(x = t \implies dx= dt\) \(y = t^2 \implies dy = 2t ~dt \)

OpenStudy (rational):

plug these in ur line integral and evaluate

OpenStudy (rational):

\[\int_C x dx + (2x+3y)dy = \int_0^1 t ~dt + (2t + 3t^2) 2tdt\]

OpenStudy (rational):

you can evaluate that

OpenStudy (anonymous):

10/3?

OpenStudy (anonymous):

find the flux of the vector field f=x,y,z> acros the slanted surface of the cone z^2=x^2+y^2 for 0<=z<=1: normal vector points in the positive z direction

OpenStudy (anonymous):

you do double integral of f*nds

OpenStudy (rational):

I need to revise flux, and yes its just a double integral

OpenStudy (anonymous):

its F n dS

OpenStudy (rational):

earlier, you found normal vector for cone : \(\left< \dfrac{x}{z}, \dfrac{y}{z} , -1 \right>\)

OpenStudy (anonymous):

yeah we use that vector but since they want positive z dont you flip your signs

OpenStudy (rational):

you mean : \(\left< -\dfrac{x}{z}, -\dfrac{y}{z} , 1 \right>\) like this ?

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

or not sure

OpenStudy (rational):

\(\overrightarrow{F} \bullet \hat{n} = \left<x,y,z\right> \bullet \dfrac{\left< -\dfrac{x}{z}, -\dfrac{y}{z} , 1 \right>}{\sqrt{2}} \)

OpenStudy (rational):

i think you're right. see if above looks okay

OpenStudy (rational):

\(\sqrt{2}\) is there becoz you want a unit vector, i have just divided the normal vector by its magnitude..

OpenStudy (anonymous):

how did you get sqrt of 2

OpenStudy (rational):

\(\sqrt{2}\) is the magnitude of "normal vector"

OpenStudy (rational):

\( \Bigg|\left< -\dfrac{x}{z}, -\dfrac{y}{z} , 1 \right>\Bigg| = \sqrt{\dfrac{x^2}{z^2} + \dfrac{y^2}{z^2} + 1^2} = \sqrt{\dfrac{x^2+y^2}{z^2} + 1} = \sqrt{\dfrac{z^2}{z^2} + 1} = \sqrt{2}\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i dont think it works i think we have to use the divergence theorem

OpenStudy (rational):



OpenStudy (rational):

we want the flux only through the slant surface right ?